OpenStudy (2yoututu11):

What is the approximate area of a circle with a radius of 28 centimeters? Use π ≈ 3.14. A. 87.9 square centimeters B. 175.8 square centimeters C. 2461.8 square centimeters D. 4923.5 square centimeters

OpenStudy (anonymous):

Same process, @2yoututu11, as I showed you earlier. \(\ \huge A = \pi r^2 \). Therefore, \(\huge A= \pi 28^2 \). That's \(\ \huge 784 \pi \). Which is approximately \(\ \huge 2461.8 cm \). Note that you use \(\ \huge 3.14 \) when multiplying, not \(\ \huge \pi \).